# Explaining Tournament Solutions with Minimal Supports

### v2

@@ Abstract
Tournaments are widely used models to represent pairwise dominance between candidates, alternatives, or teams. We study the problem of providing certified explanations for why a candidate appears among the winners under various tournament rules. To this end, we identify minimal supports—minimal sub-tournaments in which the candidate is guaranteed to win regardless of how the rest of the tournament is completed (that is, the candidate is a necessary winner of the sub-tournament). This notion corresponds to an abductive explanation for the question,“Why does the winner win the tournament”—a central concept in formal explainable AI. We focus on common tournament solutions: the top cycle, the uncovered set, the Copeland rule, the Borda rule, the maximin rule, and the weighted uncovered set. For each rule we determine the size of the smallest minimal supports, and we present polynomial-time algorithms to compute them for all but the weighted uncovered set, for which the problem is NP-complete. Finally, we show how minimal supports can serve to produce compact, certified, and intuitive explanations.
@@ 1 Introduction
Tournaments are well-known mathematical structures to represent pairwise contests between alternatives, be they a set of candidates in an election, different teams in a sport league, or competing political proposals. A number of tournament solutions have been proposed and studied in the literature to aggregate the information represented in such graphs and select a set of winning alternatives. For an introduction, see Brandt, Brill, and Harrenstein (2016) and Fischer, Hudry, and Niedermeier (2016).
@@
As simple tools for collective decision making, tournament solutions are valuable for digital democracy. However, their adoption will remain limited to small-scale, low-stakes elections until further efforts are made to make them more transparent and accessible to users  (Grossi et al. 2024). Leveraging techniques from explainable AI, in this paper we propose to supplement tournament solutions with a compact certified explanation, enabling voters to efficiently verify the winning alternative while at the same time gaining a clearer understanding of the tournament solution.
@@
Explainable AI aims at helping human users to comprehend and trust the results and output created by AI systems. To this end, multiple properties have been introduced to define effective explanations (see, e.g., Miller, 2019, for an introduction). First, an explanation should contain the right quantity of information, ensuring sufficiency without excess (see, e.g., the irreducibility and validity properties by Amgoud and Ben-Naim, 2022, and Amgoud, Cooper, and Salim, 2024). Second, to foster trust, only true statements supported by evidences should be used (Kulesza et al. 2013). Third, the information provided must be relevant to the problem at hand, and, finally, an explanation should be brief, organized and non ambiguous (see, e.g., empirical evidence by  Delaunay et al.,2025). Grice (1975) refers to these four criteria as quantity, quality, relation and manner.
@@
Since tournament solutions are white-box models, a natural candidate for an explanation is to provide voters with the full information on the vote counts, to be able to verify the solution. However, this solution necessitates more communication effort than necessary, does not provide any new insight to understand the outcome, and is only accessible to experts or at least computer-savvy people (at least for the more advanced tournament solutions).
@@
In the literature, three main options have been proposed to explain and justify the choice made by a given tournament solution. First, in social choice and welfare economics (see, e.g., Arrow, Sen, and Suzumura (2010)) one is presented with a set of desirable axiomatic properties or an intuitive mathematical structure that constitutes the normative justification of a tournament solution (Cailloux and Endriss 2016; Boixel, Endriss, and de Haan 2022; Nardi, Boixel, and Endriss 2022; Peters et al. 2020). Yet, explaining why a decision is good does not necessarily help in explaining the actual decision-making process, since multiple outcomes can be supported by different good reasons. Alternatively, voters can be presented with detailed or aggregated data on their expressed preferences, in the form of statements or features of the election, manually computed or automatically extracted (Suryanarayana, Sarne, and Kraus 2022). Although more accessible, the explanations produced in this way offer no logical guarantee and may be insufficient to justify a specific result. Finally, a recent approach by Contet, Grandi, and Mengin (2024) experiments with the use of abductive and contrastive explanations from formal explanations in machine learning (see, e.g., Shih, Choi, and Darwiche (2018); Darwiche and Hirth (2020); Ignatiev et al. (2020); Marques-Silva (2022)), introducing them to a matrix-based voting setting. However, these formal explanations often are not user-friendly in view of their size and lack of visual structure.
@@
In this paper we build on the latter proposal to introduce the concept of minimal support (MS) for tournaments, which encapsulates the core reasons why a tournament winner belongs to the winner set. We define minimal supports of a winning alternative $w$ for a given tournament solution $S$ on tournament $G$ as the set of all minimal sub-tournaments of $G$ for which $w$ is a necessary winner, i.e., $w$ is the winner of $S$ in all completions of the minimal support. MSs provide a rigorous base to explain the outcome of a tournament, since they correspond to abductive explanations to the question “Why is $w$ the $S$-winner of tournament GG?”. In order to show how our explanations can be made more accessible, we provide efficient algorithms to select for the smallest MSs (SMS) and we show how to build compact and accessible explanations in natural language using MSs as election certificates.
Tournament
Solution
Computing
a SMS
Lower bound
on SMS size
Upper bound
on SMS size
@@
Weighted
Uncovered Set
NP-complete n+m2n+m-2 (n+1)(m1)(n+1)(m-1)
Table 1: Overview of our results for various tournament solutions with $n$ voters and $m$ candidates. All bounds are tight. Closed-form expressions of the SMS size are also provided for every solution but the weighted uncovered set.
After introducing preliminary notions in Section 2, we introduce in Section 3 the concept of minimal support for tournaments. While the problem of finding abductive explanations is often intractable (in its general form the decision problem is $\Sigma_{2}^{\text{P}}$}-hard Liberatore,2005), in Section 4 we provide polynomial algorithms for computing smallest minimal supports and closed-form expression of their size for the top cycle ($\operatorname{TC}$), the uncovered set ($\operatorname{UC}$), the Copeland rule ($\operatorname{CO}$), the Borda rule ($\operatorname{BO}$), and the maximin rule ($\operatorname{MM}$). We also show that the problem is NP-complete for the weighted uncovered set ($\operatorname{wUC}$). Additionally, we bound the size of the smallest MSs and show that they share a common underlying structure. In Table 1 we provide a detailed overview of our technical results. Finally, in Section 5 we show how MSs can be used to deliver certified, compact, and intuitive explanations in natural language.
@@ 1.1 Related Work
We identified three approaches in the literature aiming to explain voting outcomes. First, in the realm of voting with rankings,Suryanarayana, Sarne, and Kraus (2022) compared experimentally crowdsourced arguments with algorithmic explanations using features of a preference profile such as various notions of score. They showed that both approaches produce comparable results in terms of acceptance and legitimacy of the process, and identified a positive effect on the acceptance of the winner by the less satisfied users. Second, a stream of papers mentioned in the introduction (Cailloux and Endriss 2016; Boixel, Endriss, and de Haan 2022; Nardi, Boixel, and Endriss 2022; Peters et al. 2020) developed a logical calculus based on axiomatic properties to justify why a candidate should be the winner in a preference profile, building on the normative justification of voting rules. Third, building on recent work in formal explainable AI, Contet, Grandi, and Mengin (2024) explored the use of abductive explanations in voting with rankings for the Borda rule.
@@
In formal explainability, abductive explanations have a dual structure called contrastive explanations (Ignatiev et al. 2020). If the former answers to questions such as “why $a$ is the result of classifier $F$ on input II?”, the latter answers questions of the form “why $b$ is not the result of classifier $F$ on input II?”. These counterfactual questions are answered by identifying minimal sets of input features that have the potential of reverting the classifier result, or that, if left unchanged, can guarantee that the same outcome is reproduced. When applied to voting, contrastive explanations are strongly related to the problem of bribery, studied under various settings ranging from entire changes of a voter’s ballot to limited modifications (see, e.g., Faliszewski, Hemaspaandra, and Hemaspaandra, 2009, and Elkind, Faliszewski, and Slinko, 2009). The setting considered in this paper, where each pairwise comparison can be individually changed ignoring the transitivity of the preferences, was introduced by Faliszewski et al. (2009).
@@
Bribery in tournaments has been used to define a notion of margin of victory by Brill, Schmidt-Kraepelin, and Suksompong (2020), later extended by Döring and Peters (2023), in an effort refine the outcome of tournament solutions. Since contrastive explanations are answers to “why not?” questions, they are great tools to show why a losing alternative was not selected. However, explaining individually why each losing alternative did not win is not necessarily a good explanation of why the winner was elected (Lipton 1990).
@@
While bribery is well-studied, its dual problem—defending against it, akin to finding abductive explanations—could have been explored with similar depth. However, to our knowledge, only Chen et al. (2021) studies the computational complexity of protection against both constructive and destructive bribery, showing that protection is generally harder than the bribery problem itself.
@@
Our approach is aligned with the work of Beynier et al. (2024), who provide explanations for the lack of local envy-free allocations in fair division using formal methods akin to abductive explanations.
@@ 2 Preliminaries
Tournaments are well-studied structures used to represent compactly pairwise comparisons among alternatives. In the following presentation we use the framework of Aziz et al. (2015). Formally, a partial tournament is a pair $G=(\mathcal{C},E)$ where $\mathcal{C}$ is a nonempty finite set of candidates or alternatives and $E\subseteq\mathcal{C}\times\mathcal{C}$ is an asymmetric relation on $\mathcal{C}$, i.e., $(y,x)\not\in E$ whenever $(x,y)\in E$. A tournament $T$ is a partial tournament $(\mathcal{C},E)$ for which $E$ is also complete, i.e., either $(x,y)\in E$ or $(y,x)\in E$ for all distinct $x,y\in\mathcal{C}$.
Let $n$ be a strictly positive integer. A partial $n$-weighted tournament is a pair $G=(\mathcal{C},\mu)$ where $\mathcal{C}$ is a nonempty finite set of candidates and $\mu:\mathcal{C}\times\mathcal{C}\to\{0,\dots,n\}$ a weight function such that for all distinct $x,y\in\mathcal{C}$, $\mu(x,y)+\mu(y,x)\leq n$. We also assume $\mu(x,x)=0$ for all candidate $x$. A (complete) $n$-weighted tournament satisfies for all distinct $x,y\in\mathcal{C}$, $\mu(x,y)+\mu(y,x)=n$. Naturally, (unweighted) tournaments can be seen as 1-weighted tournaments.
@@
Consider tournament $G=(\mathcal{C},E)$ in Figure 1(a), the arrow from node $a$ to node $b$ indicates that $a$ beats $b$ in a pairwise comparison. In the 5-weighted tournament $G_{w}=(\mathcal{C},E_{w})$ in Figure 1(b), the “3” on the arrow from $a$ to $b$ means that $a$ beats $b$ in three pairwise comparisons.
@@
In the rest of this paper, edges with a weight of zero are not represented.
@@
Let $G=(\mathcal{C},E)$ be a complete tournament. A nonempty subset of candidates $A\subseteq\mathcal{C}$ is called dominant in $G$ if for all candidate $x\in A$, for all alternative $y\in\mathcal{C}\setminus{A}$, $(x,y)\in E$. A candidate $x\in\mathcal{C}$ is said to cover another candidate $y\in\mathcal{C}$ if for all $z\in\mathcal{C}$, $(y,z)\in E\implies(x,z)\in E$.
  • the top cycle ($\operatorname{TC}$) is the (unique) minimal dominant nonempty subset of candidates of $G$. Alternatively, it can also be defined as the set of candidates that can reach every other candidate via a directed path in $G$.
  • the uncovered set ($\operatorname{UC}$) is the nonempty subset of candidates not covered by any other candidate in $G$. Alternatively, it can also be defined as the set of candidates that can reach every other candidate via a directed path in $G$ of length at most two.
  • the Copeland score of a candidate of $c\in\mathcal{C}$ in $G$ is $\sigma_{\operatorname{CO}}(c,G)=\sum_{c^{\prime}\in\mathcal{C}}\mathds{1}_{\{(c,c^{\prime})\in E\}}$. The Copeland rule ($\operatorname{CO}$) selects candidates that have a maximal Copeland score.
@@
  • the Borda score of a candidate of $c\in\mathcal{C}$ in $G$ is $\sigma_{\operatorname{BO}}(c,G)=\sum_{c^{\prime}\in\mathcal{C}}\mu(c,c^{\prime})$. The Borda rule ($\operatorname{BO}$) selects candidates that have a maximal Borda score.
  • the maximin score of a candidate $c\in\mathcal{C}$ in $G$ is $\sigma_{\operatorname{MM}}(c,G)=\min_{c^{\prime}\in\mathcal{C}\setminus{\{c\}}}\mu(c,c^{\prime})$. The maximin rule ($\operatorname{MM}$) selects candidates that have a maximal maximin score.
  • the weighted uncovered set ($\operatorname{wUC}$) is the nonempty subset of candidates not weighted covered by any other candidate in $G$. A path-based alternative definition also exists (see lemma 7).
Observe that these six tournament solutions can be organized in two families. First, there are the path-based solutions: $\operatorname{TC}$, $\operatorname{UC}$ and $\operatorname{wUC}$ which are solutions associated to some notion of paths. Second, $\operatorname{CO}$, $\operatorname{BO}$ and $\operatorname{MM}$ are rules which score candidates where each candidate’s score can be computed using only the pairwise comparisons between him and the other candidates. We call them myopic-score solutions.

### v3

@@ Abstract
Tournaments are widely used models to represent pairwise dominance between candidates, alternatives, or teams. We study the problem of providing certified explanations for why a candidate appears among the winners under various tournament rules. To this end, we identify minimal supports—minimal sub-tournaments in which the candidate is guaranteed to win regardless of how the rest of the tournament is completed (that is, the candidate is a necessary winner of the sub-tournament). This notion corresponds to an abductive explanation for the question,“Why does the winner win the tournament?”—a central concept in formal explainable AI. We focus on common tournament solutions: the top cycle, the uncovered set, the Copeland rule, the Borda rule, the maximin rule, and the weighted uncovered set. For each rule we determine the size of the smallest minimal supports, and we present polynomial-time algorithms to compute them for all solutions except for the weighted uncovered set, for which the problem is NP-complete. Finally, we show how minimal supports can serve to produce compact, certified, and intuitive explanations for tournament solutions.
@@ 1 Introduction
Tournaments are well-known mathematical structures to represent pairwise contests between alternatives, be they a set of candidates in an election, different teams in a sport league, or competing political proposals. A number of tournament solutions have been proposed and studied in the literature to aggregate the information represented in such graphs and select a set of winning alternatives. For an introduction, see tournamentHandbook2016 and weight-tournamentHandbook2016.
@@
As simple tools for collective decision making, tournament solutions are valuable for digital democracy. However, their adoption will remain limited to small-scale, low-stakes elections until further efforts are made to make them more transparent and accessible to users  (GrossiDigitalDem2024). Leveraging techniques from explainable AI (XAI), in this paper we propose to supplement tournament solutions with a compact certified explanation, enabling voters to efficiently verify the winning alternative while at the same time gaining a clearer understanding of the tournament solution.
@@
XAI aims at helping human users to comprehend and trust the outputs of AI systems. To this end, multiple properties have been introduced to define effective explanations (see, e.g., miller2019explanation miller2019explanation, for an introduction). First, an explanation should contain the right quantity of information, ensuring sufficiency without excess (see, e.g., the irreducibility and validity properties by amgoud2022axiomatic amgoud2022axiomatic, and amgoud2024axiomatic amgoud2024axiomatic). Second, to foster trust, only true statements supported by evidences should be used (kulesza2013too). Third, the information provided must be relevant to the problem at hand. Finally, an explanation should be brief, organized and non ambiguous (see, e.g., empirical evidence by  delaunay2025impact delaunay2025impact). grice1975logic refers to these four criteria as quantity, quality, relation and manner.
@@
Since tournament solutions are white-box models, a natural candidate for an explanation is to provide voters with the full information on the vote counts, to be able to recompute the solution and verify it. However, this approach requires more communication effort than necessary, does not provide any new insight to understand the outcome, and is only accessible to experts or at least computer-savvy people for the more advanced tournament solutions.
@@
In the literature, three main options have been proposed to explain and justify the choice returned by a given tournament solution. First, in social choice and welfare economics (see, e.g., arrow2010handbook arrow2010handbook) one is presented with a set of desirable axiomatic properties or an intuitive mathematical structure that constitutes the normative justification of a tournament solution (CaillouxEndrissAAMAS2016; boixel2022calculus; NardiEtAlAAMAS2022; peters2020explainable). Yet, explaining why a decision is good does not necessarily help in explaining the actual decision-making process, since multiple outcomes can be supported by different good reasons. Alternatively, voters can be presented with detailed or aggregated data on their expressed preferences, in the form of statements or features of the election, manually computed or automatically extracted (SuryanarayanaEtAlAAMAS2022). Although more accessible, the explanations produced in this way offer no logical guarantee and may be insufficient to justify a specific result. Finally, a recent approach by contet2024abductive experiments with the use of abductive and contrastive explanations from formal explanations in machine learning (see, e.g., ijcai2018p708 ijcai2018p708, darwiche2020reasons darwiche2020reasons, ignatiev_relating_2020 ignatiev_relating_2020, marques-silva_logic-based_2022 marques-silva_logic-based_2022), introducing them to a matrix-based voting setting. However, these formal explanations often are not user-friendly in view of their size and lack of visual structure.
@@
In this paper we build on the latter proposal to introduce the concept of minimal supports (MS) for tournaments solutions, which encapsulate the core reasons why a tournament winner belongs to the winner set. We define minimal supports of a winning alternative $w$ for a given tournament solution $S$ on a tournament $G$ as the set of all minimal sub-tournaments of $G$ for which $w$ is a necessary winner, i.e., $w$ is the winner of $S$ in all completions of the MS. MSs provide a rigorous base to explain the outcome of a tournament, as they correspond to abductive (or prime implicant) explanations to the question “Why is $w$ the $S$-winner of tournament GG?”.  [^1] Abduction has been defined in a variety of ways in XAI for ML and in logic-based explanations. In particular, sometimes it is limited to “why not” questions (see, e.g., eiter1995complexity eiter1995complexity). In this paper, we follow the terminology of ignatiev_relating_2020. To show how our explanations can be made accessible, we provide efficient algorithms to select the smallest MSs (SMS) and we show how to build rigorous, compact and accessible explanations in natural language using MSs.
Tournament Solution Computing an SMS Lower bound on SMS size Upper bound on SMS size
@@
Weighted Uncovered Set NP-complete n+m2n+m-2 (n+1)(m1)(n+1)(m-1)
Table 1: Overview of our results for various tournament solutions with $n$ voters and $m$ candidates. All bounds are tight. Closed-form expressions of the SMSs size are also provided for every solution but the weighted uncovered set.
After introducing preliminary notions in Section 2, we introduce in Section 3 the concept of minimal support for tournaments solutions. While the problem of finding smallest abductive explanations is often intractable (without constraints on the classifier, the decision problem is $\Sigma_{2}^{\text{P}}$}-hard (liberatore2005redundancy)), in Section 4 we provide polynomial algorithms for computing smallest minimal supports and closed-form expressions of their size for the top cycle ($\operatorname{TC}$), the uncovered set ($\operatorname{UC}$), the Copeland rule ($\operatorname{CO}$), the Borda rule ($\operatorname{BO}$), and the maximin rule ($\operatorname{MM}$). We also show that the problem is NP-complete for the weighted uncovered set ($\operatorname{wUC}$). Additionally, we bound the size of the smallest MSs and show that have underlying structure. In Table 1 we provide a detailed overview of our technical results. Finally, in Section 5 we show how MSs can be used to deliver certified, compact, and intuitive explanations in natural language.
@@ 1.1 Related Work
We identified three approaches in the literature aiming to explain voting outcomes. First, in the realm of voting with rankings, SuryanarayanaEtAlAAMAS2022 compared experimentally crowdsourced arguments with algorithmic explanations using features of a preference profile such as various notions of score. They showed that both approaches produce comparable results in terms of acceptance and legitimacy of the process, and identified a positive effect on the acceptance of the winner by the less satisfied users. Second, a stream of papers mentioned earlier (CaillouxEndrissAAMAS2016; boixel2022calculus; NardiEtAlAAMAS2022; peters2020explainable) developed a logical calculus based on axiomatic properties to justify why a candidate should be the winner in a preference profile, building on the normative justification of voting rules. Third, building on recent work in formal XAI, contet2024abductive explored the use of abductive explanations in voting with rankings for the Borda rule.
@@
In formal explainability, abductive explanations have a dual structure called contrastive explanations (ignatiev_relating_2020). If abductive explanations answer to questions such as “Why is $a$ the result of classifier $F$ on input II?”, contrastive explanations answer questions of the form “Why is $b$ not the result of classifier $F$ on input II?”. These counterfactual questions are answered by identifying minimal sets of input features that have the potential of reverting the classifier result, or that, if left unchanged, can guarantee that the same outcome is reproduced. When applied to voting, contrastive explanations are strongly related to the problem of bribery, studied under various settings ranging from entire changes of a voter’s ballot to limited modifications (see, e.g., faliszewski2009hard faliszewski2009hard, and elkind2009swap elkind2009swap). The setting considered in this paper, where each pairwise comparison can be individually changed ignoring the transitivity of the preferences, was introduced by faliszewski2009llull.
@@
The study of bribery in tournaments has lead brill2020refining to define the notion of margin of victory, later extended by doring2023margin, in an effort to refine the outcome of tournament solutions. Since contrastive explanations are answers to “why not?” questions, they are great tools to show why a losing alternative was not selected. However, explaining individually why each losing alternative did not win is not necessarily a good explanation of why the winner was elected (lipton1990contrastive).
@@
While bribery is well-studied, its dual problem—defending against it, akin to finding abductive explanations—could have been explored in similar depth. However, to our knowledge, only chen2021computational studies the computational complexity of protection against both constructive and destructive bribery, showing that protection is generally harder than the bribery problem itself.
@@
Our approach is in line with the work of beynier2024explaining, who provide explanations for the lack of local envy-free allocations in fair division using formal methods akin to abductive explanations and present them with graphs.
@@ 2 Preliminaries
Tournaments are well-studied structures used to represent compactly pairwise comparisons among alternatives. In the following presentation we use the framework of aziz2015possible. Formally, a partial tournament is a pair $G=(\mathcal{C},E)$ where $\mathcal{C}$ is a nonempty finite set of candidates or alternatives and $E\subseteq\mathcal{C}\times\mathcal{C}$ is an asymmetric relation on $\mathcal{C}$, i.e., $(y,x)\not\in E$ whenever $(x,y)\in E$. A tournament $G$ is a partial tournament $(\mathcal{C},E)$ for which $E$ is also complete, i.e., either $(x,y)\in E$ or $(y,x)\in E$ for all distinct $x,y\in\mathcal{C}$.
Let the number of voters $n$ be a strictly positive integer. A partial $n$-weighted tournament is a pair $G=(\mathcal{C},\mu)$ where $\mathcal{C}$ is a nonempty finite set of candidates and $\mu:\mathcal{C}\times\mathcal{C}\to\{0,\dots,n\}$ a weight function such that for all distinct $x,y\in\mathcal{C}$, $\mu(x,y)+\mu(y,x)\leq n$. We also assume $\mu(x,x)=0$ for every candidate $x$. A (complete) $n$-weighted tournament satisfies for all distinct $x,y\in\mathcal{C}$, $\mu(x,y)+\mu(y,x)=n$. Naturally, (unweighted) tournaments can be seen as 1-weighted tournaments.
@@
Consider tournament $G=(\mathcal{C},E)$ in Figure 1(a), the arrow from node $a$ to node $b$ indicates that $a$ is preferred to $b$, or, that $a$ beats $b$ in a pairwise comparison. In the 5-weighted tournament $G_{w}=(\mathcal{C},E_{w})$ in Figure 1(b), the “3” on the arrow from $a$ to $b$ means that 3 voters prefer $a$ to $b$, or, that $a$ beats $b$ in 3 pairwise comparisons.
@@
In the remainder of this paper, edges of weighted tournament with a weight of zero are not represented.
@@
Let $G=(\mathcal{C},E)$ be a complete tournament. A nonempty subset of candidates $A\subseteq\mathcal{C}$ is called dominant in $G$ if for every candidate $x\in A$, for all candidates $y\in\mathcal{C}\setminus{A}$, $(x,y)\in E$. A candidate $x\in\mathcal{C}$ is said to cover another candidate $y\in\mathcal{C}$ if for all $z\in\mathcal{C}$, $(y,z)\in E\implies(x,z)\in E$.
  • The top cycle ($\operatorname{TC}$) is the (unique) minimal dominant nonempty subset of candidates of $G$, i.e., the top strongly connected component of $G$. Alternatively, it can also be defined as the set of candidates that can reach every other candidate via a directed path in $G$.
  • The uncovered set ($\operatorname{UC}$) is the (unique) nonempty subset of candidates not covered by any other candidate in $G$. Alternatively, it can also be defined as the set of candidates that can reach every other candidate via a directed path in $G$ of length at most two.
  • The Copeland score of a candidate $c\in\mathcal{C}$ in $G=(\mathcal{C},E)$ is $\sigma_{\operatorname{CO}}(c,G)=|\{c^{\prime}:c^{\prime}\in\mathcal{C},(c,c^{\prime})\in E\}|$. The Copeland rule ($\operatorname{CO}$) selects candidates that have a maximal Copeland score.
@@
  • The Borda score of a candidate $c\in\mathcal{C}$ in $G=(\mathcal{C},\mu)$ is $\sigma_{\operatorname{BO}}(c,G)=\sum_{c^{\prime}\in\mathcal{C}}\mu(c,c^{\prime})$. The Borda rule ($\operatorname{BO}$) selects candidates that have a maximal Borda score.
  • The maximin score of a candidate $c\in\mathcal{C}$ in $G=(\mathcal{C},\mu)$ is $\sigma_{\operatorname{MM}}(c,G)=\min_{c^{\prime}\in\mathcal{C}\setminus{\{c\}}}\mu(c,c^{\prime})$. The maximin rule ($\operatorname{MM}$) selects candidates that have a maximal maximin score.
  • The weighted uncovered set ($\operatorname{wUC}$) is the nonempty subset of candidates that are not weighted covered by any other candidate in $G$. A path-based alternative definition also exists (see Lemma 7).
Observe that these six tournament solutions can be organized in two families. First, there are the path-based solutions: $\operatorname{TC}$, $\operatorname{UC}$ and $\operatorname{wUC}$ which are solutions associated to some notion of paths. Second, $\operatorname{CO}$, $\operatorname{BO}$ and $\operatorname{MM}$ are rules which score candidates where each candidate’s score can be computed using only the pairwise comparisons between it and the other candidates. We call them myopic-score solutions. This distinction will become useful in Section 3.3 when discriminating between MSs.